# Boson star-black hole binaries: initial data and head-on collisions

### 16 Apr 2026 v1

@@ I Introduction
Several DM candidates have been proposed, ranging from particle-like candidates such as weakly interacting massive particles (WIMPs) Pospelov et al. (2008); Arcadi et al. (2018); Roszkowski et al. (2018); Schumann (2019) and sterile neutrinos Dodelson and Widrow (1994); Shi and Fuller (1999); Abazajian and others (2012), to wave-like candidates such as ultralight scalar fields Hu et al. (2000); Hui et al. (2017); Hui (2021), and even macroscopic objects such as primordial black holes (PBHs) Carr et al. (2016); Green and Kavanagh (2021); Ning et al. (2026b); Zeng et al. (2025); Carr et al. (2026); Ning et al. (2026a); Klipfel et al. (2026); Jia et al. (2026). Among these candidates, scalar fields are particularly well motivated and have been studied extensively. They arise naturally in a wide range of theoretical scenarios, including the QCD axion Dine et al. (1981); Dine and Fischler (1983); Preskill et al. (1983), axion-like particles Marsh (2016) from string-theory compactifications Svrcek and Witten (2006); Arvanitaki et al. (2010), and more general scalar-field DM models. If endowed with mass, scalar fields minimally coupled to gravity can form macroscopic self-gravitating structures Kaup (1968); Ruffini and Bonazzola (1969); Seidel and Suen (1991). Complex scalar fields can support static, spherically symmetric configurations in which the field oscillates in time, known as boson stars (BSs) Kaup (1968); Ruffini and Bonazzola (1969), while real scalar fields can instead form long-lived oscillating configurations with a nontrivial time-dependent energy-momentum tensor, called oscillatons Seidel and Suen (1991) (see, e.g., Refs. Schunck and Mielke (2003); Liebling and Palenzuela (2023); Visinelli (2021); Bezares and Sanchis-Gual (2025) for reviews). Both types of solutions can arise naturally as end states of gravitational collapse Kaup (1968); Ruffini and Bonazzola (1969); Seidel and Suen (1991); Okawa et al. (2014a); Chavanis (2016); Widdicombe et al. (2018) and have a maximum mass scaling as $M_{\mathrm{max}}\sim M_{\mathrm{pl}}^{2}/\mu$, where $M_{\mathrm{pl}}$ is the Planck mass and $\mu$ is the scalar-field mass. Consequently, depending on the scalar mass, these objects can span a wide range of masses and play very different cosmological roles.
@@
At the same time, axion and axion-like scenarios, as well as string-inspired “axiverse” constructions, motivate a much broader mass range of scalar fields. For heavier scalars, such as $\mu\sim 10^{-10}-10^{-16}\,\mathrm{eV}$, the associated self-gravitating configurations can span masses from solar to supermassive scales. In addition, modifying the scalar self-interaction potential can further enlarge the allowed ranges of mass and compactness, enabling configurations that behave as exotic compact objects (ECOs) or, in some regimes, as black-hole (BH) mimickers Cardoso et al. (2014); Marks et al. (2025); Staelens (2026); Evstafyeva et al. (2026). The mergers of such objects can produce strong GW signals that may be detectable by current and future GW observatories, and these signals may be distinguishable from standard BH or neutron-star (NS) mergers through their characteristic features.
@@
Motivated by these properties, BS-BS binary mergers have been extensively investigated in numerical relativity, including head-on collisions Bernal and Siddhartha Guzman (2006); Palenzuela et al. (2007); Choptuik and Pretorius (2010); Mundim (2010); Cardoso et al. (2016); Helfer et al. (2022); Jaramillo et al. (2022); Sanchis-Gual et al. (2022b); Evstafyeva et al. (2023b); Atteneder et al. (2024); Siemonsen (2024); Lira et al. (2025); Ge et al. (2025); Brito et al. (2026); Ge (2025) and orbital inspirals Palenzuela et al. (2008); Mundim (2010); Palenzuela et al. (2017); Bezares et al. (2017); Bezares and Palenzuela (2018); Sanchis-Gual et al. (2020); Bezares et al. (2022); Croft et al. (2023); Siemonsen and East (2023a, b) for both equal- Palenzuela et al. (2007, 2008); Choptuik and Pretorius (2010); Mundim (2010); Cardoso et al. (2016); Palenzuela et al. (2017); Bezares et al. (2017); Bezares and Palenzuela (2018); Sanchis-Gual et al. (2020); Helfer et al. (2022); Jaramillo et al. (2022); Sanchis-Gual et al. (2022b); Croft et al. (2023); Siemonsen and East (2023a); Atteneder et al. (2024); Lira et al. (2025); Ge et al. (2025); Brito et al. (2026); Ge (2025) and unequal-mass Bernal and Siddhartha Guzman (2006); Bezares et al. (2022); Evstafyeva et al. (2023b); Siemonsen and East (2023b); Siemonsen (2024) binaries. These studies have revealed rich phenomenology, including waveform degeneracies with BH binaries and the imprint of tidal deformability. Furthermore, other ECO mergers have also been studied, such as oscillaton Helfer et al. (2019); Widdicombe et al. (2020) and Proca-star (vector cousins of BSs) Sanchis-Gual et al. (2019); Calderón Bustillo et al. (2021); Calderon Bustillo et al. (2023); Sanchis-Gual et al. (2022a) binaries. From an astrophysical perspective, if BSs, fuzzy-DM cores, or other scalar-field configuratons populate the Universe, their coexistence with BHs and their inevitable interactions are highly probable. In this context, the nonlinear dynamics of scalar fields coupled to BHs—including scalar-field accretion Sanchis-Gual et al. (2015); Hui et al. (2019); Clough et al. (2019); Cardoso et al. (2022a); Brito et al. (2015a, 2016), scalarization Okawa et al. (2014b); Silva et al. (2018); Zhang et al. (2022); Doneva et al. (2024); Chen et al. (2023b); Guo et al. (2026), descalarization Silva et al. (2021); Chen et al. (2023a), superradiance Brito et al. (2015b); Baryakhtar et al. (2021); Lyu et al. (2025), dynamical friction Traykova et al. (2021); Boudon et al. (2022), and dynamical transition processes Ning et al. (2024); Chen et al. (2023c); Zhao et al. (2026b, a)—have been widely explored in various strongly gravitating systems, revealing rich time-evolution behaviors and phase structures. It is therefore also crucial to understand the dynamics and GW signatures of mixed BS-BH binary mergers for interpreting future observations and constraining the existence of such objects.
@@ II.1 Model
where $g$ is the determinant of the spacetime metric $g_{\mu\nu}$, $R$ and $\nabla_{\mu}$ denotes the Ricci scalar and the covariant derivative associated with this metric, $\bar{\varphi}$ is the complex conjugate of the scalar field, and $V(|\varphi|^{2})$ is the scalar-field potential. Varying the action with respect to the metric and the scalar field yields the Einstein-Klein-Gordon equations:
@@ II.3 Diagnostics
We also extract the outgoing GW radiation in terms of the Newman-Penrose scalar $\Psi_{4}$ Newman and Penrose (1962); Bishop and Rezzolla (2016); Radia et al. (2022) and decompose it into spin-weight s=2s=-2 spherical harmonics $Y^{-2}_{\ell m}$ Bruegmann et al. (2008) according to
@@
We also compute the cumulative GW energy $E_{\mathrm{GW}}$ from the Newman-Penrose scalar multipoles $\Psi_{4}^{\ell m}$ via
$\dot{E}_{\mathrm{GW}}^{\ell m}(t)$ $=\frac{R_{\mathrm{ex}}^{2}}{16\pi}\left|\int_{-\infty}^{t}\Psi_{4}^{\ell m}\mathrm{d}\tilde{t}\right|^{2},$ (18)
$\dot{E}_{\mathrm{GW}}(t)$ $=\sum_{\ell,m}\dot{E}_{\mathrm{GW}}^{\ell m}(t).$ (19)
We begin the time integration at $t=50$ to exclude any spurious or “junk” radiation present in the initial data.
@@ III.2 Lorentz boost
The extrinsic curvature are computed analytically from $\tilde{K}_{\tilde{i}\tilde{j}}=-\frac{1}{2\tilde{\alpha}}(\partial_{\tilde{t}}-\mathcal{L}_{\tilde{\beta}})\tilde{\gamma}_{\tilde{i}\tilde{j}}$ Croft (2022). The nonzero components are
@@
$\quad\left.-\frac{\tilde{x}\cosh\eta}{\tilde{r}}A^{\prime}(\sinh\eta+\tilde{\beta}^{\tilde{x}}\cosh\eta)\right]e^{-i\omega\tilde{x}\sinh\eta}.$
@@ III.3 Plain superposition and its malaise
away from the isolated equilibrium value $\gamma_{ij}^{\mathrm{A}}(x_{\mathrm{A}}^{i})$. Physically, this perturbation distorts the volume element $\sqrt{\gamma}$ at the BS core, effectively injecting spurious mass-energy and compressing the BS away from its delicate equilibrium. In practice, this artificial compression increases the Hamiltonian-constraint violation, induces unphysical radial oscillations, alters the GW signal, and, for sufficiently compact models (especially solitonic BSs), can trigger premature collapse into a BH before the physical merger begins. This is the principal “malaise” of plain superposition in BS-BH initial data: even when the asymptotic behavior is acceptable, the local BS core can already be driven away from equilibrium at $t=0$.
@@ IV Numerical results
For the BS-BH binaries, we consider mass ratios $q=M_{\mathrm{BS}}/M_{\mathrm{BH}}=0.38$, $0.5$, 11, 22, and $2.64$, and we also include several BS-BS and BH-BH cases for comparison. To test the validity of the improved superposition method, we additionally include a plain-superposition case for the equal-mass BS-BH binary and vary the initial separation to assess the severity of the artifacts. The parameters of the binary configurations are listed in Table 2. The initial velocities are chosen so that the total initial momentum vanishes, while the initial relative velocity is kept fixed at $\Delta v=v_{A}-v_{B}=0.2$ for all cases. For all BS-BH binaries, we use the improved superposition method described in Sec. III.4. For BSBH-q1-H, we also include the plain superposition method to illustrate the severity of the artifacts it produces. For the BS-BS binaries, we use the improved superposition method of Ref. Evstafyeva et al. (2023b) for all cases. We uniformly set the correction-radius parameter to $R_{\mathrm{BS}}=50$ for all BSs, which works well for the configurations considered here. For the BH-BH binaries, we use the plain superposition method, which is known to be sufficient for BH-BH initial data. For runs with $M_{\mathrm{BH}}=0.2700$, we use an additional refinement level to resolve the smaller BH.
@@ IV.1.1 Initial constraint violations
Figure 1: Comparison of the initial Hamiltonian (top panel) and momentum (bottom panel) constraint violations along the collision axis for the configuration BSBH-q1-H with $d=20$, generated using the plain (solid lines) and improved (dashed lines) superposition methods, respectively. The improved superposition method significantly reduces the Hamiltonian-constraint violation in the vicinity of both the BS (at $x=246$) and the BH (at $x=266$) relative to the plain superposition method.
@@ IV.3 Unequal-mass BS-BH collisions
Before examining the GW emission from unequal-mass binaries, we must address a major distinction between BS-BS and BS-BH systems in their initial parameter space. As emphasized by Evstafyeva et al. Evstafyeva et al. (2023b), the merger dynamics and radiated energy of unequal-mass BS-BS binaries are highly sensitive to the relative scalar phase at the moment of collision. Because unequal-mass BSs have different characteristic frequencies ($\omega_{A}\neq\omega_{B}$), they inevitably accumulate a dynamical phase difference $\Delta\phi\approx(\omega_{A}-\omega_{B})t$ during the infall. Consequently, setting the initial phase difference to zero at $t=0$ leads to a merger phase—and therefore to constructive or destructive interference patterns—that depends strongly on the initial separation $d$. This strong phase degeneracy makes it difficult to define a representative GW energy for unequal-mass BS-BS collisions without an exhaustive scan over the phase space.
@@
By contrast, the BH in a BS-BH binary carries no scalar phase. The interaction and accretion dynamics are governed solely by the BS stress-energy tensor, which is invariant under global $U(1)$ phase transformations. The merger dynamics of BS-BH systems are therefore intrinsically deterministic, robust, and independent of any phase offset. To ensure a rigorous and unambiguous comparison across binary parameters, we exclude the strongly phase-dependent BS-BS configurations from the following analysis of unequal-mass systems. Instead, we focus on contrasting the phase-independent BS-BH collisions with their pure BH-BH counterparts across different mass ratios $q$.
@@
We first examine the radiated GW energy, shown in the top panels of Fig. 6 for mass ratios $q\in\{0.5,2\}$ (left) and $q\in\{0.38,2.64\}$ (right). When the BH is the heavier companion ($q<1$, orange solid lines), the radiative efficiency of the BS-BH binary is consistently lower than that of the corresponding BH-BH system (blue dotted lines), as in the equal-mass case. However, a more intricate nonlinear competition emerges when the highly compact BS is the heavier object ($q>1$, green solid lines). For $q=2$, the radiated energy of the BS-BH system slightly exceeds that of the pure BH-BH collision. This is consistent with the recent findings of Marks et al. Marks et al. (2026), who observed that sufficiently compact BS-BH collisions can radiate more efficiently than their BH-BH counterparts in the $q>1$ regime. Interestingly, when the mass ratio is increased further to $q=2.64$, the BS-BH energy drops slightly below the BH-BH baseline again. This behavior appears to contrast with Ref. Marks et al. (2026), where the positive excess in radiated energy for $q>1$ was found to grow with $q$ up to $q=3$. We attribute this difference to the distinct kinematic setups. Marks et al. adopted a constant initial velocity ($v=0.1$) for all objects, thereby assigning an artificially large kinetic energy to the more massive BS. In contrast, we use a rigorously controlled center-of-momentum frame ($P_{x}=0$, $\Delta v=0.2$), which is more physically realistic. In the $q=2.64$ configuration, the much lighter BH must move with a relatively large initial velocity ($v_{\mathrm{BH}}=0.1447$) to plunge into the massive, slowly moving BS ($v_{\mathrm{BS}}=0.0553$). This “bullet-like” piercing induces strong dynamical friction, transferring a substantial fraction of the BH kinetic energy into internal scalar excitations and scalar wakes rather than into quadrupolar GWs. Thus, in $q>1$ mixed binaries, the radiated GW energy reflects a delicate competition between the increased emission associated with the larger mass ratio and the dissipative effects of kinematic piercing, which suppress the radiation.

### 30 Apr 2026 v2

@@ I Introduction
Several DM candidates have been proposed, ranging from particle-like candidates such as weakly interacting massive particles (WIMPs) Pospelov et al. (2008); Arcadi et al. (2018); Roszkowski et al. (2018); Schumann (2019) and sterile neutrinos Dodelson and Widrow (1994); Shi and Fuller (1999); Abazajian and others (2012), to wave-like candidates such as ultralight scalar fields Hu et al. (2000); Hui et al. (2017); Hui (2021), and even macroscopic objects such as primordial black holes (PBHs) Carr et al. (2016); Green and Kavanagh (2021); Ning et al. (2026b); Zeng et al. (2025); Ning et al. (2025); Carr et al. (2026); Ning et al. (2026a); Klipfel et al. (2026); Jia et al. (2026). Among these candidates, scalar fields are particularly well motivated and have been studied extensively. They arise naturally in a wide range of theoretical scenarios, including the QCD axion Dine et al. (1981); Dine and Fischler (1983); Preskill et al. (1983), axion-like particles Marsh (2016) from string-theory compactifications Svrcek and Witten (2006); Arvanitaki et al. (2010), and more general scalar-field DM models. If endowed with mass, scalar fields minimally coupled to gravity can form macroscopic self-gravitating structures Kaup (1968); Ruffini and Bonazzola (1969); Seidel and Suen (1991). Complex scalar fields can support static, spherically symmetric configurations in which the field oscillates in time, known as boson stars (BSs) Kaup (1968); Ruffini and Bonazzola (1969), while real scalar fields can instead form long-lived oscillating configurations with a nontrivial time-dependent energy-momentum tensor, called oscillatons Seidel and Suen (1991) (see, e.g., Refs. Schunck and Mielke (2003); Liebling and Palenzuela (2023); Visinelli (2021); Bezares and Sanchis-Gual (2025) for reviews). Both types of solutions can arise naturally as end states of gravitational collapse Kaup (1968); Ruffini and Bonazzola (1969); Khlopov et al. (1985); Seidel and Suen (1991); Okawa et al. (2014a); Chavanis (2016); Widdicombe et al. (2018) and have a maximum mass scaling as $M_{\mathrm{max}}\sim M_{\mathrm{pl}}^{2}/\mu$, where $M_{\mathrm{pl}}$ is the Planck mass and $\mu$ is the scalar-field mass. Consequently, depending on the scalar mass, these objects can span a wide range of masses and play very different cosmological roles.
@@
At the same time, axion and axion-like scenarios, as well as string-inspired “axiverse” constructions, motivate a much broader mass range of scalar fields. For heavier scalars, such as $\mu\sim 10^{-10}-10^{-16}\,\mathrm{eV}$, the associated self-gravitating configurations can span masses from solar to supermassive scales. In addition, modifying the scalar self-interaction potential can further enlarge the allowed ranges of mass and compactness, enabling configurations that behave as exotic compact objects (ECOs) or, in some regimes, as black-hole (BH) mimickers Cardoso et al. (2014); Marks et al. (2025); Staelens (2026); Evstafyeva et al. (2026); Wang et al. (2026). The mergers of such objects can produce strong GW signals that may be detectable by current and future GW observatories, and these signals may be distinguishable from standard BH or neutron-star (NS) mergers through their characteristic features.
@@
Motivated by these properties, BS-BS binary mergers have been extensively investigated in numerical relativity, including head-on collisions Bernal and Siddhartha Guzman (2006); Palenzuela et al. (2007); Choptuik and Pretorius (2010); Mundim (2010); Cardoso et al. (2016); Helfer et al. (2022); Jaramillo et al. (2022); Sanchis-Gual et al. (2022b); Evstafyeva et al. (2023b); Atteneder et al. (2024); Siemonsen (2024); Lira et al. (2025); Ge et al. (2025); Brito et al. (2026); Ge (2025) and orbital inspirals Palenzuela et al. (2008); Mundim (2010); Palenzuela et al. (2017); Bezares et al. (2017); Bezares and Palenzuela (2018); Sanchis-Gual et al. (2020); Bezares et al. (2022); Croft et al. (2023); Siemonsen and East (2023a, b) for both equal- Palenzuela et al. (2007, 2008); Choptuik and Pretorius (2010); Mundim (2010); Cardoso et al. (2016); Palenzuela et al. (2017); Bezares et al. (2017); Bezares and Palenzuela (2018); Sanchis-Gual et al. (2020); Helfer et al. (2022); Jaramillo et al. (2022); Sanchis-Gual et al. (2022b); Croft et al. (2023); Siemonsen and East (2023a); Atteneder et al. (2024); Lira et al. (2025); Ge et al. (2025); Brito et al. (2026); Ge (2025) and unequal-mass Bernal and Siddhartha Guzman (2006); Bezares et al. (2022); Evstafyeva et al. (2023b); Siemonsen and East (2023b); Siemonsen (2024) binaries. These studies have revealed rich phenomenology, including waveform degeneracies with BH binaries and the imprint of tidal deformability. Furthermore, other ECO mergers have also been studied, such as oscillaton Helfer et al. (2019); Widdicombe et al. (2020) and Proca-star (vector cousins of BSs) Sanchis-Gual et al. (2019); Calderón Bustillo et al. (2021); Calderon Bustillo et al. (2023); Sanchis-Gual et al. (2022a) binaries. From an astrophysical perspective, if BSs, fuzzy-DM cores, or other scalar-field configurations populate the Universe, their coexistence with BHs and their inevitable interactions are highly probable. In this context, the nonlinear dynamics of scalar fields coupled to BHs—including scalar-field accretion Sanchis-Gual et al. (2015); Hui et al. (2019); Clough et al. (2019); Cardoso et al. (2022a); Brito et al. (2015a, 2016), scalarization Okawa et al. (2014b); Silva et al. (2018); Zhang et al. (2022); Doneva et al. (2024); Chen et al. (2023b); Guo et al. (2026), descalarization Silva et al. (2021); Chen et al. (2023a), superradiance Brito et al. (2015b); Baryakhtar et al. (2021); Lyu et al. (2025), dynamical friction Traykova et al. (2021); Boudon et al. (2022), and dynamical transition processes Ning et al. (2024); Chen et al. (2023c); Zhao et al. (2026b, a)—have been widely explored in various strongly gravitating systems, revealing rich time-evolution behaviors and phase structures. It is therefore also crucial to understand the dynamics and GW signatures of mixed BS-BH binary mergers for interpreting future observations and constraining the existence of such objects.
@@ II.1 Model
where $g$ is the determinant of the spacetime metric $g_{\mu\nu}$, $R$ and $\nabla_{\mu}$ denote the Ricci scalar and the covariant derivative associated with this metric, $\bar{\varphi}$ is the complex conjugate of the scalar field, and $V(|\varphi|^{2})$ is the scalar-field potential. Varying the action with respect to the metric and the scalar field yields the Einstein-Klein-Gordon equations:
@@ II.3 Diagnostics
We extract the outgoing GW radiation in terms of the Newman-Penrose scalar $\Psi_{4}$ Newman and Penrose (1962); Bishop and Rezzolla (2016); Radia et al. (2022) and decompose it into spin-weight s=2s=-2 spherical harmonics $Y^{-2}_{\ell m}$ Bruegmann et al. (2008) according to
@@
We also compute the cumulative GW energy $E_{\mathrm{GW}}$ from the Newman-Penrose scalar multipoles $\Psi_{4}^{\ell m}$ via Isaacson (1968)
$\dot{E}_{\mathrm{GW}}^{\ell m}(t)$ $=\frac{R_{\mathrm{ex}}^{2}}{16\pi}\left|\int_{-\infty}^{t}\Psi_{4}^{\ell m}(\tilde{t})\mathrm{d}\tilde{t}\right|^{2},$ (18)
$\dot{E}_{\mathrm{GW}}(t)$ $=\sum_{\ell,m}\dot{E}_{\mathrm{GW}}^{\ell m}(t),$ (19)
$E_{\mathrm{GW}}(t)$ $=\int_{t_{0}}^{t}\dot{E}_{\mathrm{GW}}(\tilde{t})\mathrm{d}\tilde{t}$. (20)
We begin the time integration at $t_{0}=50$ to exclude any spurious or “junk” radiation present in the initial data.
@@ III.2 Lorentz boost
The extrinsic curvature is computed analytically from $\tilde{K}_{\tilde{i}\tilde{j}}=-\frac{1}{2\tilde{\alpha}}(\partial_{\tilde{t}}-\mathcal{L}_{\tilde{\beta}})\tilde{\gamma}_{\tilde{i}\tilde{j}}$ Croft (2022). The nonzero components are
@@
$\quad\left.-\frac{x}{r}A^{\prime}(\sinh\eta+\tilde{\beta}^{\tilde{x}}\cosh\eta)\right]e^{-i\omega\tilde{x}\sinh\eta}.$
@@ III.3 Plain superposition and its malaise
away from the isolated equilibrium value $\gamma_{ij}^{\mathrm{A}}(x_{\mathrm{A}}^{i})$. Physically, this perturbation distorts the volume element $\sqrt{\gamma}$ at the BS core, effectively injecting spurious mass-energy and compressing the BS away from its delicate equilibrium. In practice, this artificial compression increases the Hamiltonian-constraint violation, induces unphysical radial oscillations, and alters the GW signal. For sufficiently compact models (especially solitonic BSs), it can even trigger premature collapse into a BH before the physical merger begins. This is the principal “malaise” of plain superposition in BS-BH initial data: even when the asymptotic behavior is acceptable, the local BS core can already be driven away from equilibrium at $t=0$.
@@ IV Numerical results
For the BS-BH binaries, we consider mass ratios $q=M_{\mathrm{BS}}/M_{\mathrm{BH}}=0.38$, $0.5$, 11, 22, and $2.64$, and we also include several BS-BS and BH-BH cases for comparison. To test the validity of the improved superposition method, we additionally include a plain-superposition case for the equal-mass BS-BH binary and vary the initial separation to assess the severity of the artifacts. The parameters of the binary configurations are listed in Table 2. The initial velocities are chosen so that the total initial momentum vanishes, while the initial relative velocity is kept fixed at $\Delta v=v_{A}-v_{B}=0.2$ for all cases. For all BS-BH binaries, we use the improved superposition method described in Sec. III.4. For BSBH-q1-H, we also include the plain superposition method to illustrate the severity of the artifacts it produces. For the BS-BS binaries, we use the improved superposition method of Ref. Evstafyeva et al. (2023b) for all cases. We uniformly set the correction-radius parameter to $R_{\mathrm{BS}}=50$ for all BSs, which works well for the configurations considered here. For the BH-BH binaries, we use the plain-superposition construction as a reference baseline; for the separations and velocities considered here, this provides a sufficient comparison dataset for the purposes of this work. For runs with $M_{\mathrm{BH}}=0.2700$, we use an additional refinement level to resolve the smaller BH.
@@ IV.1.1 Initial constraint violations
Figure 1: Comparison of the initial Hamiltonian (left panel) and momentum (right panel) constraint violations along the collision axis for the configuration BSBH-q1-H with $d=20$, generated using the plain (solid lines) and improved (dashed lines) superposition methods, respectively. The improved superposition method significantly reduces the Hamiltonian-constraint violation in the vicinity of both the BS (at $x=246$) and the BH (at $x=266$) relative to the plain superposition method.
@@ IV.3 Unequal-mass BS-BH collisions
Before examining the GW emission from unequal-mass binaries, we must address a major distinction between BS-BS and BS-BH systems in their initial parameter space. As emphasized by Evstafyeva et al. Evstafyeva et al. (2023b), the merger dynamics and radiated energy of unequal-mass BS-BS binaries are highly sensitive to the relative scalar phase at the moment of collision. Because unequal-mass BSs have different characteristic frequencies ($\omega_{A}\neq\omega_{B}$), they inevitably accumulate a dynamical phase difference $\Delta\phi\approx(\omega_{A}-\omega_{B})t$ during the infall. Consequently, setting the initial phase difference to zero at $t=0$ leads to a merger phase—and therefore to constructive or destructive interference patterns—that depends strongly on the initial separation $d$. This strong phase dependency makes it difficult to define a representative GW energy for unequal-mass BS-BS collisions without an exhaustive scan over the phase space.
@@
By contrast, the BH in a BS-BH binary carries no scalar phase. The interaction and accretion dynamics are governed solely by the BS stress-energy tensor, which is invariant under global $U(1)$ phase transformations. The merger dynamics of BS-BH systems are therefore free from any scalar-phase dependency. To ensure a rigorous and unambiguous comparison across binary parameters, we exclude the strongly phase-dependent BS-BS configurations from the following analysis of unequal-mass systems. Instead, we focus on contrasting the phase-independent BS-BH collisions with their pure BH-BH counterparts across different mass ratios $q$.
@@
We first examine the radiated GW energy, shown in the top panels of Fig. 6 for mass ratios $q\in\{0.5,2\}$ (left) and $q\in\{0.38,2.64\}$ (right). When the BH is the heavier companion ($q<1$, blue solid lines), the radiative efficiency of the BS-BH binary is consistently lower than that of the corresponding BH-BH system (blue dotted lines), as in the equal-mass case. However, a more intricate nonlinear competition emerges when the highly compact BS is the heavier object ($q>1$, orange solid lines). For $q=2$, the radiated energy of the BS-BH system slightly exceeds that of the pure BH-BH collision. This is consistent with the recent findings of Marks et al. Marks et al. (2026), who observed that sufficiently compact BS-BH collisions can radiate more efficiently than their BH-BH counterparts in the $q>1$ regime. Interestingly, when the mass ratio is increased further to $q=2.64$, the BS-BH energy drops slightly below the BH-BH baseline again. This behavior appears to contrast with Ref. Marks et al. (2026), where the positive excess in radiated energy for $q>1$ was found to grow with $q$ up to $q=3$. We attribute this difference to the distinct kinematic setups. Marks et al. adopted a constant initial velocity ($v=0.1$) for all objects, thereby assigning an artificially large kinetic energy to the more massive BS. In contrast, we use a rigorously controlled center-of-momentum frame ($P_{x}=0$, $\Delta v=0.2$), which is more physically realistic. In the $q=2.64$ configuration, the much lighter BH must move with a relatively large initial velocity ($v_{\mathrm{BH}}=0.1447$) to plunge into the massive, slowly moving BS ($v_{\mathrm{BS}}=0.0553$). This “bullet-like” piercing induces strong dynamical friction, transferring a substantial fraction of the BH kinetic energy into internal scalar excitations and scalar wakes rather than into quadrupolar GWs. Thus, in $q>1$ mixed binaries, the radiated GW energy reflects a delicate competition between the increased emission associated with the larger mass ratio and the dissipative effects of kinematic piercing, which suppress the radiation.